OpenStudy (anonymous):

The product of the page numbers on two facing pages of a book is 240. Find the page numbers.

OpenStudy (chaise):

(n*(n+1)=240 n^2+n = 240 n^2 + n - 240 = 0 n = 15 n = 16 The two pages are 15 and 16.

OpenStudy (anonymous):

Thank you again! That was my last question.